# SAT Encoding of Partial Ordering Models for Graph Coloring Problems

### v1

@@ Abstract
In this paper, we suggest new SAT encodings of the partial-ordering based ILP model for the graph coloring problem (GCP) and the bandwidth coloring problem (BCP). The GCP asks for the minimum number of colors that can be assigned to the vertices of a given graph such that each two adjacent vertices get different colors. The BCP is a generalization, where each edge has a weight that enforces a minimal ?distance? between the assigned colors, and the goal is to minimize the ?largest? color used.
For the widely studied GCP, we experimentally compare our new SAT encoding to the state-of-the-art approaches on the DIMACS benchmark set. Our evaluation confirms that this SAT encoding is effective for sparse graphs and even outperforms the state-of-the-art on some DIMACS instances.
@@ 1 Introduction
SAT approaches are based on encoding the problem as a Boolean Satisfiability problem. A possible encoding consists of introducing color variables $x_{v,i}$, where a true assignment of $x_{v,i}$ represents assigning vertex $v$ with color $i$ (e.g., [10, 3]). Other methods are based on Zykov’s tree induced by Zykov’s deletion-contration recurrence (e.g., [8, 7]), in which the models contain variables $s_{u,v}$ that encode if vertices $u$ and $v$ have the same or different colors. Heule, Karahalios and van Hoeve [10] have introduced the algorithm CliColCom in which they alternatingly solve a maximum clique problem and a graph coloring problem using SAT approaches, where the solution from one problem helps finding a solution for the other problem and vice versa.
The most natural ILP model is the assignment-based ILP model, which directly assigns colors to vertices by introducing binary variables $x_{v,i}$ that (similar to the color variables in SAT) decide if vertex $v$ is assigned to color $i$. A drawback of this formulation is the presence of symmetries in the solution space: For a valid coloring, any permutation of the color labels provides another equivalent solution leading to a significant larger search space. Mendez-Diaz and Zabala [18] have suggested additional symmetry-breaking constraints, that completely eliminate this type of symmetry. Mutzel and Jabrayilov [12] have proposed ILP formulations, which are based on formulating the coloring problem as a partial-ordering problem (POP). This model suggests ordering the colors and placing the vertices relatively in this order. It contains binary variables $y_{v,i}$ for every vertex $v$ and possible color $i$, indicating if color $i$ is smaller than vertex $v$. It has been shown that for sparse graphs, the simple assignment and partial-ordering based models show good performance [12] and that the partial-ordering based models dominate the assignment-based models. A rather complex but also competitive ILP model is based on the set covering formulation [17] (see, e.g. [9]), which uses the fact that a coloring describes a partitioning of the vertices into independent sets, and contains a variable $x_{s}$ for every independent set $s$ in the graph, that decides if $s$ is a set of vertices corresponding to a color class in the coloring. Because there can be an exponential amount of independent sets, the formulation cannot be solved using standard techniques and instead has to be solved using column generation methods.
@@
Our contribution. Motivated by the observation that methods based on SAT encodings as well as partial-ordering based ILP models have shown to be successful, we suggest SAT encodings of the partial-ordering based model for both coloring variants. For the GCP, we also strengthen the model using the symmetry-breaking constraints by Mendez-Diaz and Zabala [18] in order to eliminate the inherent symmetries in the solution space. Our experimental evaluation for the GCP shows that the new SAT encoding of the POP model outperforms the assignment-based SAT encoding as well as all evaluated ILP formulations from the literature on the DIMACS benchmark set.
@@ 2.1.1 The assignment models (ASS-I) and (ASS-I-B)
where $\bar{d}$ is the average edge distance in $G$. The first equality can be derived as follows: For every color $i$, the interval of colors $j$ satisfying $|i-j|<d(e)$ is $j\in[i-d(e)+1,i+d(e)-1]$. This interval contains exactly $12d(e)-$ 1 elements, which for the $H$ colors $i=1,...,H$ leads to $H\cdot(2d(e)-1)$ pairs $(i,j)$ in total. However, we have to subtract the pairs we counted for which $j<1$ or $j>H$. For every $i$ with $i-d(e)<1$, there exist exactly $d(e)-i+1$ pairs $(i,j)$ for which $j<1$: $(i,0),(i,-1),...,(i,i-d(e))$. In total, we have $\sum_{i=1}^{d(e)}(d(e)-i+1)=(d(e)^{2}-d(e))/2$ of such pairs. For $j>H$ the situation is symmetrical, leading to a total number of $d(e)^{2}-d(e)$ of pairs we need to subtract for each edge.
@@ 2.2 SAT encodings (ASS-S) and (ASS-S-B)
3 New formulations based on the partial-ordering approach
@@ 3.1 SAT encodings based on partial-ordering: (POP-S) and (POPH-S)
Here, we suggest new SAT encodings for the GCP which are based on the partial-ordering based model.
@@ 3.1.1 Hybrid partial-ordering based SAT encoding for the GCP
$\neg x_{v,i}\lor y_{v,i-1}$ $\forall v\in V,i=2,...,H$ (16c)
$\neg x_{v,i}\lor\neg y_{v,i}$ $\forall v\in V,i=2,...,H$ (16d)
$x_{v,i}\lor\neg y_{v,i-1}\lor y_{v,i}$ $\forall v\in V,i=2,...,H$ (16e)
@@ 4 Experimental evaluation
In our computational experiments we evaluated the effectiveness of our new SAT encodings and compared them with state-of-the art approaches. In particular, we were interested in a comparison of the partial-ordering based encoding with the assignment based SAT encoding (i.e., the basic SAT encoding) as well as the ILP formulations of the assignment and the partial-ordering based models. Moreover, we compared our new models to state-of-the-art approaches. The implementation and the data is publically available on https://github.com/s6dafabe/popsatgcpbcp.
@@ 4.2 Test setup and benchmark set
To solve the SAT encodings, we used the solver kissat 3.1.1111https://github.com/arminbiere/kissat/releases/tag/rel-3.1.1, accessed 11.03.2024, which has been successful in the 2022 SAT competition. The preprocessing and the generation of the SAT and ILP formulations were implemented in Python 3.10 using the library networkx 2.8.5222https://networkx.org/documentation/stable/release/release_2.8.5.html. For solving the ILP models, we used Gurobi 10.0.2 single-threadedly. The machine used to evaluate the SAT and ILP formulations features an Intel Xeon Gold 6130@2.1GHz running CentOS Linux and 187 GB of memory (Benchmarks [1] user time: r500.5=4.87s). We also compiled the implementation of Held, Cook and Sewell [9] with the same Gurobi version. Because their implementation [^3] https://github.com/heldstephan/exactcolors, accessed 11.03.2024 uses features of Gurobi that are incompatible with the first machine, this method had to be evaluated on a different machine having an AMD EPYC 7543P@2.8GHz and 257GB memory (Benchmarks [1] user time: r500.5=3.24s). For the graph coloring problem, we performed our experiments on a set of 134 DIMACS graphs [20] and additionally a set of 9 randomly generated instances by Michael Trick (the R-instances: Note that there exist 18 instances in total, however the instances are duplicated and the duplicates only differ in the node weights, which are irrelevant for standard graph coloring). Furthermore, we compare with the results reported in [10] of the methods presented in [10] and [7]. For similar reasons as reported in Heule, Karahalios and van Hoeve [10], we did not compare to the work in [6], which claims strong results for graph coloring with a method using a relaxed Zykov encoding that is incrementally strengthened. The linked source code is currently incorrect, and the authors were unable to reproduce the results. For the bandwidth coloring problem, we used the GEOM set consisting of 33 graphs generated by Michael Trick [20]. We have set a time limit of 1 hour.
set ASS-S POP-S POPH-S ASS-I POP-I POPH-I EC[9] CDCL[7, 10] CLICOL [10]
DSJ 2/15 2/15 2/15 3/14 2/14 3/14 5/15 2/15 3/15
FullIns 14/14 14/14 14/14 12/14 11/14 12/14 5/14 14/14 14/14
Insertions 4/11 4/11 4/11 4/11 4/11 4/11 1/11 3/11 4/11
abb313GPI 1/1 1/1 1/1 0/1 0/1 0/1 0/1 0/1 0/1
anna 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1
ash 3/3 3/3 3/3 3/3 3/3 3/3 0/3 3/3 3/3
david 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1
flat 0/6 0/6 0/6 0/6 0/6 0/6 2/6 0/6 0/6
fpsol2 3/3 3/3 3/3 3/3 1/3 3/3 3/3 3/3 3/3
games120 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1
homer 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1
huck 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1
inithx 3/3 3/3 3/3 3/3 1/3 3/3 3/3 3/3 3/3
jean 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1
latin_square 0/1 0/1 0/1 0/0 0/0 0/0 0/1 0/1 0/1
le450 8/12 8/12 8/12 8/12 7/12 8/12 3/12 8/12 9/12
miles 5/5 5/5 5/5 5/5 5/5 5/5 5/5 5/5 5/5
mug 4/4 4/4 4/4 4/4 4/4 4/4 4/4 4/4 4/4
mulsol 5/5 5/5 5/5 5/5 4/5 5/5 5/5 5/5 5/5
myciel 4/5 4/5 4/5 3/5 3/5 3/5 2/5 4/5 4/5
r 8/9 7/9 7/9 7/9 5/9 7/9 7/9 7/9 7/9
R 7/9 7/9 7/9 7/9 5/9 7/9 9/9 0/0 0/0
qg 3/4 3/4 3/4 1/3 0/3 2/3 0/4 3/4 2/4
queen 6/13 8/13 8/13 6/13 5/13 6/13 7/13 6/13 5/13
school1 2/2 2/2 2/2 2/2 2/2 2/2 1/2 1/1 1/1
wap0 3/8 5/8 5/8 0/8 0/8 1/8 1/8 1/8 4/8 444the instance wap08a was solved by the method in seperate experiments using a much higher time limit
will199GPI 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1 1/1
zeroin 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3 3/3
total 95/143 98/143 98/143 86/140 72/140 88/140 73/143 83/134 87/134
Table 1: Number of solved instances on the benchmark set for the GCP
@@ 4.3 Experimental results for the graph coloring problem
Table 1 shows the number of solved instances for the 143 evaluated instances for the SAT encodings of the assignment (ASS-S), the partial-ordering (POP-S), and the hybrid partial-ordering (POPH-S) based models, the corresponding ILP formulations (ASS-I), (POP-I), (POPH-I), the method by Held, Cook, and Sewel [9], the results of the method of Hebrard and Katsirelos [7], and those of Heule, Karahalios and van Hoeve [10] as reported in [10]. The first column of the table describes the class type and the subsequent columns show the number of solved instances for each model, out of the total number of tested instances. Note that the authors in [10] did not evaluate on the R instances, which is why the total number of evaluated instances for this model is only 134 (Note that they evaluated on 137 instances in total, however, we omitted the three C instances in our evaluations as they are very large and not solved by any method). For the ILP models, we omitted graphs with more than 300.000 edges, as this resulted in out-of-memory exceptions, resulting in 140 evaluated instances. The bold items in the table highlight the instance types for which the POP encodings solved more instances than the other methods.
We can see that (POP-S) and (POPH-S) solved the most instances (98/143). Note that even when excluding the instances which were not used in the paper by Heule, Karahalios and van Hoeve [10], (POP-S) and (POPH-S) still solved more instances (91/125) than their approach. Furthermore, the POP based SAT encodings were able to solve three more instances than (ASS-S), and there was only one instance that was solved by (ASS-S) but not by the POP based SAT encodings. The POP encodings performed especially well on the wap0-class, as they solved the instances wap01a, wap02a and wap08a, which were only closed recently [10].
@@
Figure 1 visualizes for each model the number of instances, which can be solved within a time limit of 1, 2,…, 3600 seconds. We omitted the R instances in this figure for better comparability. We can see that (POP-S) and (POPH-S) solve the most instances independent of the considered time limit. An interesting observation is that for ILP formulations, the POP formulation performs far worse than the other two formulations (ASS-I) and (POPH-I), while for the SAT encodings, (POP-S) and (POPH-S) show almost identical performance (with (POP-S) even being slightly better). Jabrayilov and Mutzel [12] argued that one weakness of the POP ILP formulaton lies in the denser constraint matrix, which is caused by the POP model containing 4 variables in the constraints enforcing differing colors for connected vertices. However, this does not seem to impact the performance of the SAT encoding.

### v2

@@ Abstract
In this paper, we revisit SAT encodings of the partial-ordering based ILP model for the graph coloring problem (GCP) and suggest a generalization for the bandwidth coloring problem (BCP). The GCP asks for the minimum number of colors that can be assigned to the vertices of a given graph such that each two adjacent vertices get different colors. The BCP is a generalization, where each edge has a weight that enforces a minimal ?distance? between the assigned colors, and the goal is to minimize the ?largest? color used.
For the widely studied GCP, we experimentally compare the partial-ordering based SAT encoding to the state-of-the-art approaches on the DIMACS benchmark set. Our evaluation confirms that this SAT encoding is effective for sparse graphs and even outperforms the state-of-the-art on some DIMACS instances.
@@ 1 Introduction
SAT approaches are based on encoding the problem as a Boolean Satisfiability problem. A possible encoding consists of introducing color variables $x_{v,i}$, where a true assignment of $x_{v,i}$ represents assigning vertex $v$ with color $i$ (e.g., [11, 4]). Other methods are based on Zykov’s tree induced by Zykov’s deletion-contration recurrence (e.g., [9, 8]), in which the models contain variables $s_{u,v}$ that encode if vertices $u$ and $v$ have the same or different colors. Heule, Karahalios and van Hoeve [11] have introduced the algorithm CliColCom in which they alternatingly solve a maximum clique problem and a graph coloring problem using SAT approaches, where the solution from one problem helps finding a solution for the other problem and vice versa. Most relevant to this work are the SAT encodings suggested by Tamura et al. [21] and Ansótegui et al. [1], which contains binary variables $y_{v,i}$ for every vertex $v$ and possible color $i$, indicating if color $i$ is smaller than vertex $v$. Although the experimental evaluation in [21] has shown that this encoding has dominated the assignment SAT encoding, these results have caught little attention in the recent literature.
The most natural ILP model is the assignment-based ILP model, which directly assigns colors to vertices by introducing binary variables $x_{v,i}$ that (similar to the color variables in SAT) decide if vertex $v$ is assigned to color $i$. A drawback of this formulation is the presence of symmetries in the solution space: For a valid coloring, any permutation of the color labels provides another equivalent solution leading to a significant larger search space. Mendez-Diaz and Zabala [19] have suggested additional symmetry-breaking constraints, that completely eliminate this type of symmetry. Mutzel and Jabrayilov [13] have proposed ILP formulations, which are based on formulating the coloring problem as a partial-ordering problem (POP). This model suggests ordering the colors and placing the vertices relatively in this order, analogous to the ordering encoding in [21] and [1] for SAT. It has been shown that for sparse graphs, the simple assignment and partial-ordering based models show good performance [13] and that the partial-ordering based models dominate the assignment-based models. Furthermore, [14] have shown theoretical advantages of POP over the assignment ILP. A rather complex but also competitive ILP model is based on the set covering formulation [18] (see, e.g. [10]), which uses the fact that a coloring describes a partitioning of the vertices into independent sets, and contains a variable $x_{s}$ for every independent set $s$ in the graph, that decides if $s$ is a set of vertices corresponding to a color class in the coloring. Because there can be an exponential amount of independent sets, the formulation cannot be solved using standard techniques and instead has to be solved using column generation methods.
@@
Our contribution. Motivated by the recent interest of the ILP community in partial-ordering based ILP models, we revisited SAT encodings of the partial-ordering based model for the GCP and generalize them to the BCP. For the GCP, we also strengthen the model using the symmetry-breaking constraints by Mendez-Diaz and Zabala [19] in order to eliminate the inherent symmetries in the solution space. Our experimental evaluation for the GCP shows that the partial-ordering based SAT encoding of the POP model outperforms the assignment-based SAT encoding as well as all evaluated ILP formulations from the literature on the DIMACS benchmark set.
@@ 2.1.1 The assignment models (ASS-I) and (ASS-I-B)
where $\bar{d}$ is the average edge distance in $G$. The first equality can be derived as follows: For every color $i$, the interval of colors $j$ satisfying $|i-j|<d(e)$ is $j\in[i-d(e)+1,i+d(e)-1]$. This interval contains exactly $12d(e)-$ 1 elements, which for the $H$ colors $i=1,...,H$ leads to $H\cdot(2d(e)-1)$ pairs $(i,j)$ in total. However, we have to subtract the pairs we counted for which $j<1$ or $j>H$. For every $i$ with $i-d(e)<1$, there exist exactly $d(e)-i$ pairs $(i,j)$ for which $j<1$: $(i,0),(i,-1),...,(i,i-d(e)+1)$. In total, we have $\sum_{i=1}^{d(e)}(d(e)-i)=(d(e)^{2}-d(e))/2$ of such pairs. For $j>H$ the situation is symmetrical, leading to a total number of $d(e)^{2}-d(e)$ of pairs we need to subtract for each edge.
@@ 2.2 SAT encodings (ASS-S) and (ASS-S-B)
3 Formulations based on the partial-ordering approach
Here, we revisit the SAT encoding suggested by [21] and [1] for the GCP which also can be seen as the SAT-counterpart to the partial-ordering based ILP model. We suggest a modification of the symmetry-breaking constraints used in [19] for the partial-ordering based model that can be encoded into SAT in polynomial size and without adding new variables. Furthermore, we suggest a new hybrid version inspired by the (POPH-I) model.
@@ 3.1.1 Hybrid partial-ordering based SAT encoding for the GCP
$\neg x_{v,i}\lor y_{v,i-1}$ $\forall v\in V,i=2,...,k$ (16c)
$\neg x_{v,i}\lor\neg y_{v,i}$ $\forall v\in V,i=2,...,k$ (16d)
$x_{v,i}\lor\neg y_{v,i-1}\lor y_{v,i}$ $\forall v\in V,i=2,...,k$ (16e)
@@ 4 Experimental evaluation
In our computational experiments, we evaluated the effectiveness of the partial-ordering based SAT encodings and compared them with state-of-the art approaches. In particular, we were interested in a comparison of the partial-ordering based encoding with the assignment based SAT encoding (i.e., the basic SAT encoding) as well as the ILP formulations of the assignment and the partial-ordering based models. Moreover, we compared the models to state-of-the-art approaches. The implementation and the data is publically available on https://github.com/s6dafabe/popsatgcpbcp.
@@ 4.2 Test setup and benchmark set
To solve the SAT encodings, we used the solver kissat 3.1.1111https://github.com/arminbiere/kissat/releases/tag/rel-3.1.1, accessed 11.03.2024, which has been successful in the 2022 SAT competition. The preprocessing and the generation of the SAT and ILP formulations were implemented in Python 3.10 using the library networkx 2.8.5222https://networkx.org/documentation/stable/release/release_2.8.5.html. For solving the ILP models, we used Gurobi 10.0.2 single-threadedly. The machine used to evaluate the SAT and ILP formulations features an Intel Xeon Gold 6130@2.1GHz running CentOS Linux and 187 GB of memory (Benchmarks [2] user time: r500.5=4.87s). For comparison, we compiled the implementation of Heule, Karahalios and Hoeve [11] and tested it on the same machine. We also compiled the implementation of Held, Cook and Sewell [10] with the same Gurobi version. Because their implementation [^3] https://github.com/heldstephan/exactcolors, accessed 11.03.2024 uses features of Gurobi that are incompatible with the first machine, this method had to be evaluated on a different machine having an AMD EPYC 7543P@2.8GHz and 257GB memory (Benchmarks [2] user time: r500.5=3.24s). For the graph coloring problem, we performed our experiments on a set of 134 DIMACS graphs [22] and additionally a set of 9 randomly generated instances by Michael Trick (the R-instances: Note that there exist 18 instances in total, however the instances are duplicated and the duplicates only differ in the node weights, which are irrelevant for standard graph coloring). Furthermore, we compare with the results reported in [11] of the method presented in [8]. For similar reasons as reported in Heule, Karahalios and van Hoeve [11], we did not compare to the work in [7], which claims strong results for graph coloring with a method using a relaxed Zykov encoding that is incrementally strengthened. The linked source code is currently incorrect, and the authors were unable to reproduce the results. For the bandwidth coloring problem, we used the GEOM set consisting of 33 graphs generated by Michael Trick [22]. We have set a time limit of 1 hour.
@@ 4.3 Experimental results for the graph coloring problem
set ASS-S POP-S POPH-S ASS-I POP-I POPH-I EC[10] CLICOL [11] CDCL[8, 11]  [^4] The code provided in the repository produced compile errors on our system, so we used the results of the experiments from [11] which did not contain the R-instances
DSJ 2 2 2 3 2 3 5 4 2
FullIns 14 14 14 12 11 12 5 14 14
Insertions 4 4 4 4 4 4 1 4 3
abb313GPI 1 1 1 0 0 0 0 0 0
anna 1 1 1 1 1 1 1 1 1
ash 3 3 3 3 3 3 0 3 3
david 1 1 1 1 1 1 1 1 1
flat 0 0 0 0 0 0 2 0 0
fpsol2 3 3 3 3 1 3 3 3 3
games120 1 1 1 1 1 1 1 1 1
homer 1 1 1 1 1 1 1 1 1
huck 1 1 1 1 1 1 1 1 1
inithx 3 3 3 3 1 3 3 3 3
jean 1 1 1 1 1 1 1 1 1
latin_square 0 0 0 0 0 0 0 0 0
le450 8 8 8 8 7 8 3 10 8
miles 5 5 5 5 5 5 5 5 5
mug 4 4 4 4 4 4 4 4 4
mulsol 5 5 5 5 4 5 5 5 5
myciel 4 4 4 3 3 3 2 4 4
r 8 7 7 7 5 7 7 7 7
qg 3 3 3 1 0 2 0 3 3
queen 6 8 8 6 5 6 7 5 6
school1 2 2 2 2 2 2 1 2 1
wap0 3 5 5 0 0 1 1 4 1
will199GPI 1 1 1 1 1 1 1 1 1
zeroin 3 3 3 3 3 3 3 3 3
total 88 91 91 79 67 81 64 90 83
Table 1: Number of solved DIMACS instances on the benchmark set for the GCP
set ASS-S POP-S POPH-S ASS-I POP-I POPH-I EC[10] CLICOL [11]
R50 3 3 3 3 3 3 3 3
R75 2 2 2 2 1 2 3 3
R100 2 2 2 2 1 2 3 2
total 7 7 7 7 5 7 9 8
Table 2: Number of solved instances on the R-instances
Table 1 shows the number of solved instances for the 134 evaluated DIMACS instances for the SAT encodings of the assignment (ASS-S), the partial-ordering (POP-S), and the hybrid partial-ordering (POPH-S) based models, the corresponding ILP formulations (ASS-I), (POP-I), (POPH-I), the method by Held, Cook, and Sewel [10] (EC), the method of Heule, Karahalios and van Hoeve [11] (CLICOL) and the results of the method of Hebrard and Katsirelos [8] (CDCL) as reported in [11]. The first column of the table describes the class type and the subsequent columns show the number of solved instances for each model, out of the total number of tested instances. Table 2 shows the number of solved instances for the 9 R-instances. Unfortunatly, the code provided  [8] did not compile on our machine, therefore we used the results of the experiments performed in [11] for the algorithm CDCL. Note that the authors in [11] did not evaluate on the R instances, which is why they are missing in table 2. Note that for the ILP models, the large instances DSJC1000.9, latin_square_10 and qg.order100 resulted in out-of-memory exceptions. Also, the algorithm CLICOL does not seem to be robust for large instances, as it produced runtime errors for the instances r1000.5, latin_square_10 and wap04a. The bold items in the table highlight the instance types for which the POP encodings solved more instances than the other methods.
We can see that (POP-S), (POPH-S) solved the most DIMACS instances (91/134), followed closely by CLICOL (90/134). Furthermore, the POP based SAT encodings were able to solve three more instances than (ASS-S), and there was only one instance that was solved by (ASS-S) but not by the POP based SAT encodings. The POP encodings performed especially well on the wap0-class, as they solved the instances wap01a, wap02a and wap08a, which were only closed recently [11].
For the R-instances, we can see that all four SAT methods perform similarly, with CLICOL slightly outperforming the three simple encodings. (ASS-I) and (POPH-I) also solve the same amount of instances as the three evaluated SAT encodings. EC performs the best, solving all instances, while (POP-I) performs the worst, which is to be expected, as the instance set contains many instances with medium and high density, for which EC typically performs well and (POP-I) typically performs poorly.
@@
Figure 1 visualizes for each model the number of instances, which can be solved within a time limit of 1, 2,…, 3600 seconds. We omitted the R instances in this figure for better comparability. We can see that (POP-S) and (POPH-S) solve more instances than (ASS-S) independent of the considered time limit. Generally, (POP-S), (POPH-S) and CLICOL are the best approaches and perform similarly. An interesting observation is that for the ILP formulations, the POP formulation performs far worse than the other two formulations (ASS-I) and (POPH-I), while for the SAT encodings, (POP-S) and (POPH-S) show almost identical performance (with (POP-S) even being slightly better). Jabrayilov and Mutzel [13] argued that one weakness of the POP ILP formulaton lies in the denser constraint matrix, which is caused by the POP model containing 4 variables in the constraints enforcing differing colors for connected vertices. However, this does not seem to impact the performance of the SAT encoding.
In total (for DIMACS and R-instances combined), (POP-S), (POPH-S) and CLICOL solved the most instances (98/143). One interesting thing to note is that although (POP-S)/(POPH-S) and CLICOL solved the same number of instances, they solved a different set of instances. For example (POP-S)/(POPH-S) is particularly advantageous on the queen instances, while CLICOL shows superior performance on the le450 instances. We want to remark that the CLICOL approach uses the assignment-based SAT encoding as a sub-algorithm and combines it with a more sophisticated method of finding an initial clique used for precoloring. An interesting idea could be to use the partial-ordering based SAT encoding in the CLICOL framework to try and combine the advantages of both methods.